# Fractional phase jumps in stochastic systems with tilted periodic double-well potentials

### v1

@@ III.2.3 Retrapping
It is often difficult to reach the steady state in experimental realizations or realistic Langevin simulations. This is especially problematic for systems with low noise and weak damping. However, once we have the rates of the most relevant processes, we can investigate even long-time processes via the deterministic Eq. (35).
As an illustration, we analyze the scenario of a sudden quench of the parameters. We employ the master equation (35) and also large-scale Langevin simulations. For the latter, we used the modified Euler-Heun method LambaEulerHeun from the Julia package DifferentialEquations.jl with adaptive integration step. In the simulations, we used $\approx 6.5\times 10^{4}$ particles and, as before, their immediate regime was identified via the separatrices for the model parameters after the quench.
@@
The calculated time dependence of the occupation probabilities are shown in Fig. 8 by the solid (master equation) and dashed (Langevin dynamics) black ($P_{R}$), blue ($P_{A}$) and red ($P_{B}$) lines. The logarithmic time scale reveals several metastable regimes. At short times, there are the largest differences between the simplified model and the full simulation, as expected. In particular, the change in initial occupations is faster for the simplified model than in the simulation. This is due to the details of the initial state, that is, the initial distribution of the particles. Full dynamics starts with a relatively high mean velocity of the particles, and therefore it takes a while to slow them down.
@@
First, particles are quickly caught in the minima around $\tau\sim 10^{1}$, because the retrapping rates are $\Gamma_{rA}\sim\Gamma_{rB}\approx 10^{-1}$. The system is close to be completely trapped before $\tau\sim 10^{2}$. However, although relatively stable, the occupation probabilities of the locked states are far from their steady-state values (horizontal dashed-dotted lines) obtained by the MCF method discussed above. This metastable state exists due to the large separation of the retrapping rates from the escape and phase jumps ones, see Fig. 7(c). Probabilities start to approach the true steady state only for $\tau\gtrsim 10^{4}$ where first the escape processes with rates $\Gamma_{eA}\sim\Gamma_{eB}\sim 10^{-4}$ and then phase jumps with $\Gamma_{AB}\sim 0.5\times 10^{-4}$ start to be relevant.
@@ IV Conclusion
For the strong damping, the analysis of the velocivy Fano factor revealed a region where single-phase jumps, including fractional ones, are the prevailing source of the velocity noise. We have shown that with decreasing temperature, the steady-state properties of the overdamped junction approach an effective single well system for any finite $\alpha$.
In the intermediate and weak damping regime, the FCS analyzes showed complex phase dynamics related to the single and multiple phase jumps. The revealed large differences between the rates of odd (fractional) and even phase jumps can be explained by analyzing the particle trajectories. Even in this regime we have identified parameter ranges, for which a single-harmonic analysis is sufficient for the description of the main phase slip statistics.
In the analysis of the switching regime, we have presented a simple master equation method to calculate the escape and retrapping rates. We have focused on both the regime where the double well character plays an important role in the steady state statistics and the regime where it does not. We have shown how this property is related to the retrapping, escape, and phase jump rates. We have demonstrated how these rates evolve with the bias force and determined the probabilities of steady-state occupations. The important observation is that retrapping, escape, and phase jumps can have rates that differ by orders of magnitude. This sets distinct time scales in realistic retrapping processes.
@@
To wrap it up, besides providing a simple technique for analyzing the statistical properties of stochastic systems with double-well periodic potentials and analyzing their properties, we have shown two rather general features of such systems that can be crucial for the analysis of experimental setups. First, for a broad range of parameters in a strong and intermediate damping regime, the steady-state system can be approached with a simple single-harmonic model. Consequently, a two-well character of real potential can be hidden in the averaged data when the escape rate from one of the minima significantly exceeds the other rates. This can be true even for wells of similar depths at low noise.
@@
Acknowledgments.— M.Ž. and T.N. acknowledge support by the Czech Science Foundation through Project No. 23-05263K. This work was supported by the Ministry of Education, Youth and Sports of the Czech Republic through the e-INFRA CZ (ID:90254). E.G. thanks Deutsche Forschungsgemeinschaft (DFG project GO-1106/6) for financial support of this collaboration. W.B. was financially supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) via the SFB 1432 (project ID 425217212). M.Ž. thanks Róbert Jurčík for rewriting the code for stochastic dynamics into Julia.

### v2

@@ III.2.3 Retrapping
It is often difficult to reach the steady state in experimental realizations or realistic Langevin simulations. This is especially problematic for systems with low noise and weak damping. However, once we have the rates of the most relevant processes, we can investigate even long-time processes via the master equation (34).
As an illustration, we analyze the scenario of a sudden quench of the parameters. We employ the master equation (34) as well as large-scale Langevin simulations. For the latter, we used the modified Euler-Heun method LambaEulerHeun from the Julia package DifferentialEquations.jl with adaptive integration step. In the simulations, we used $\approx 6.5\times 10^{4}$ particles and, as before, their immediate regime was identified via the separatrices for the model parameters after the quench.
@@
The calculated time dependencies of the occupation probabilities are shown in Fig. 9 by the solid (master equation) and dashed (Langevin dynamics) black ($P_{R}$), blue ($P_{A}$) and red ($P_{B}$) lines. The logarithmic time scale reveals several metastable regimes. At short times, there are the largest differences between the simplified model and the full simulation, as expected. In particular, the change in initial occupations is faster for the simplified model than in the simulation. This is due to the details of the initial state, that is, the initial distribution of the particles. Full dynamics starts with a relatively high mean velocity of the particles, and therefore it takes a while to slow them down.
@@
First, particles are quickly caught in the minima around $\tau\sim 10^{1}$, because the retrapping rates are $\Gamma_{rA}\sim\Gamma_{rB}\approx 10^{-1}$. The system is close to being completely trapped before $\tau\sim 10^{2}$. However, although relatively stable, the occupation probabilities of the locked states are far from their steady-state values (horizontal dash-dotted lines) obtained by the MCF method discussed above. This metastable state exists due to the large separation of the retrapping rates from the escape and positional jumps rates; see Fig. 7(c). Probabilities start to approach the true steady state only for $\tau\gtrsim 10^{4}$ where first the escape processes with rates $\Gamma_{eA}\sim\Gamma_{eB}\sim 10^{-4}$ and then positional jumps with $\Gamma_{AB}\sim 0.5\times 10^{-4}$ start to be relevant.
@@ IV Conclusion
For strong damping, analysis of the velocity Fano factor revealed a region where single jumps, including fractional ones, are the prevailing source of the velocity noise. We have shown that with decreasing temperature, the steady-state properties of the overdamped junction approach an effective single well system for any finite $\alpha$.
In the intermediate and weak damping regime, the FCS analysis showed complex dynamics related to the single and multiple positional jumps. The revealed large differences between the rates of odd (fractional) and even positional jumps can be explained by analyzing the particle trajectories. Even in this regime, we have identified parameter ranges, for which a single-harmonic analysis is sufficient for the description of the main positional slip statistics.
In the analysis of the switching regime, we have presented a simple master equation method to calculate the escape and retrapping rates. We have focused on both the regime where the double-well character plays an important role in the steady state statistics and the regime where it does not. We have shown how this property is related to the retrapping, escape, and positional jump rates. We have demonstrated how these rates evolve with the bias force and determined the probabilities of steady-state occupations. The important observation is that retrapping, escape, and positional jumps can have rates that differ by orders of magnitude. This sets distinct time scales in realistic retrapping processes.
@@
To wrap up, besides providing a simple technique for analyzing the statistical properties of stochastic systems with double-well periodic potentials and analyzing their properties, we have shown two rather general features of such systems that can be crucial for the analysis of experimental setups. First, for a broad range of parameters in a strong and intermediate damping regime, the steady-state system can be approached with a simple single-harmonic model. Consequently, the two-well character of real potentials can be hidden in the averaged data when the escape rate from one of the minima significantly exceeds the other rates. This can be true even for wells of similar depths at low noise.
@@
Acknowledgments.— M.Ž. and T.N. acknowledge support by the Czech Science Foundation through Project No. 23-05263K. This work was supported by the Ministry of Education, Youth and Sports of the Czech Republic through the e-INFRA CZ (ID:90254). E.G. thanks Deutsche Forschungsgemeinschaft (DFG Project No. GO-1106/6) for financial support of this collaboration. W.B. was financially supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) via the SFB 1432 (Project ID 425217212). M.Ž. thanks Róbert Jurčík for rewriting the code for stochastic dynamics into Julia.